OpenStudy (anonymous):

What is the value of m in the following equation, 10m - (7m+4) - 5m = 16? -10 -6 1 5/2

OpenStudy (anonymous):

10m - (7m+4) - 5m = 16 10m - 7m - 4 - 5m = 16 -2m - 4 = 16 12 = 2m 6 = m

OpenStudy (anonymous):

you mean 12=-2m m=-6

OpenStudy (anonymous):

look at your equations again, you put -2m-4=16 add 4 to both sides and you get -2m=12, you dropped the negative sign on 2m for no reason

OpenStudy (anonymous):

YEs! m=-6 Sorry.